Please help answer the following questions on how operating systems handle memory allocation in Main Memory.

  1. Suppose page size = 2^11 and the logical address is 15 bits. How many pages can a process have at most?

  2. Convert the logical address of 18723 to a <page#, offset> format. Assume the page size from the question 1.

  3. Suppose the logical address 18723 is located in frame 5. What is the physical address that the logical address is mapped to?
Answer #1

Answer:

i)W have given,

Logical address = 15 bits

We know that logical address = number of pages + offset

Page size = 2^11, so number of pages = 4

Logical address = number of pages + offset

15 = 4 + offest

Therefore offset = 11

So the pages processed = 4 pages

ii) Logical address we have given = 18723

Let we convert it into binary

18723 = 1001001001000011

We know logical address = pages + offset

Pages = 4 bits

Offset = 11 bits

iii)Logical address= 1001 + 001001000011 , 5 = frame

So total frame number = 101001000011 = 10819